OpenStudy (anonymous):

Solve the following system of equations by linear combination: 2d + e = 8 d e = 4 (Points : 4) The solution is (5, 2). The solution is (4, 0). There is no solution. There are an infinite number of solutions.

OpenStudy (anonymous):

(2d+e)+(d-e) = 8+4, which is 3d=12. Solve this for d, then use that value of d to solve for e.